Hello pls help me solve.
The sum of two numbers is equal to 31. The smaller is doubled. The larger is decreased by 11. The resulting numbers are equal. What is the number?

Answer by stanbon(75887)

The sum of two numbers is equal to 31. The smaller is doubled. The larger is decreased by 11. The resulting numbers are equal. What is the number?
-------
Equations:
x + y = 31
(2x)= (y-11)
---------------
Substitute for "x" and solve for "y"::
(2(31-y)) = y-11
----
62-2y = y-11
3y = 73
------
y = 26 1/2
----
x = 31- 26 1/2 = 4.5
----------------
Cheers,
Stan H.
----------